OpenStudy (anonymous):

Find the inverse Laplace transform of the given function:

OpenStudy (anonymous):

\[\huge F(s)=\frac{2(s-1)e^{-2s}}{s^2-2s+2}\] I got as far as to complete the square in the denominator so that \[F(s)=\frac{2(s-1)e^{-2s}}{(s-1)^2+1}\]I can also see that we're going to end up with a cos somewhere given the general form of F(s)..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Would it be: \[\large 2e^t \cos(t) u_2(t)\] or\[\large 2e^{t-2} \cos(t-2)u_2(t)\]

OpenStudy (irishboy123):

sort out the algebra, (s-1)^2

ganeshie8 (ganeshie8):

\[\mathcal{L}(u_a(t)f(t-a)) = e^{-as}F(s)\] \[\mathcal{L(e^{at}f(t))} = F(s-a)\]

OpenStudy (astrophysics):

\[F(s)=\frac{2(s-1)e^{-2s}}{(s-1)^2+1} = 2\mathcal{L} ^{-1}\left( e^{-2s} \times \frac{ (s-1) }{ (s-1)^2+1 } \right)\] then use \[\mathcal{L}^{-1}(e^{ct}f(t)) = F(t-c)\]

OpenStudy (astrophysics):

I like to split it like that, as it's clearer I think

OpenStudy (astrophysics):

Mine? That's how our prof does it haha

ganeshie8 (ganeshie8):

no, mine

OpenStudy (astrophysics):

Well it's just one more extra step

OpenStudy (astrophysics):

you have \[2u_2(t)f(t-2)\] then you can find f which is \[e^tcos(t)\] plug it all in \[2u_2 (t)e^{t-2}cos(t-2)\] actually I'm not entirely sure if we put t-2 in the exponent

OpenStudy (astrophysics):

I actually just learnt all this about 10 minutes ago doing a problem rofl

OpenStudy (anonymous):

Here's what I did... \[G(s)=\frac{2(s-1)}{(s-1)^2+1} \rightarrow \mathcal{L}^{-1}G(s)=2e^t \cos(t)\] Then because of the e^(-2s), then t becomes t-2 \[2e^{t-2} \cos(t-2)u_2(t)\]

OpenStudy (astrophysics):

Yeah that looks good to me, but confirm with @ganeshie8

ganeshie8 (ganeshie8):

confirm with wolfram http://www.wolframalpha.com/input/?i=inverse+laplace+%5Cfrac%7B2%28s-1%29e%5E%7B-2s%7D%7D%7B%28s-1%29%5E2%2B1%7D look at alternate forms

OpenStudy (astrophysics):

What's with the thet?

OpenStudy (astrophysics):

theta*

OpenStudy (astrophysics):

Oh ok that's there way of saying u2(t)?

ganeshie8 (ganeshie8):

the same unit step funciton, \(\theta(t-a) = u_a(t) \)

OpenStudy (anonymous):

Yeah -- I don't understand those alternate forms..

OpenStudy (astrophysics):

I gotcha

OpenStudy (irishboy123):

Wolf uses theta as the heaviside/ step function

OpenStudy (astrophysics):

cos(2-t) mhm

ganeshie8 (ganeshie8):

@CShrix \(u_a(t)\) is a shortcut for \(u(t-a)\)

ganeshie8 (ganeshie8):

you get \(u_a(t)\) by translating the unit step function to right by \(a\) units

OpenStudy (anonymous):

Ahh gotcha.. Yeah why is it cos(2-t)? Shouldn't it be t-2?

ganeshie8 (ganeshie8):

does it matter

OpenStudy (anonymous):

I think it does matter.. wouldn't that be a huge difference between -t and t?

OpenStudy (anonymous):

The exponent has t-2, but the cos has 2-t..?

ganeshie8 (ganeshie8):

recall, cos is symmetric about y axis : \[\cos(x) = \cos(-x)\]

OpenStudy (astrophysics):

Pretty much cosx is an even function where sinx is odd

OpenStudy (anonymous):

So even the sign of the 2 doesn't matter either? cos(2-x)=cos(x-2)

OpenStudy (anonymous):

Grr haha

ganeshie8 (ganeshie8):

\[2-x = -(x-2)\] happens... :P

OpenStudy (anonymous):

asdfasdf lol

OpenStudy (astrophysics):

http://www.wolframalpha.com/input/?i=cos%282-x%29%3Dcos%28x-2%29 just to confirm

OpenStudy (anonymous):

XD Gotcha now.. Thanks everyone! X)

OpenStudy (astrophysics):

Thanks for the question :D good practice hehe

OpenStudy (anonymous):

You're welcome X) I have an exam on Tuesday and I haven't exactly had a lot of time recently to stay ahead of the game.. :(

OpenStudy (astrophysics):

Ah darn, my final is about 2 weeks away lol

OpenStudy (anonymous):

Same.. We have an exam on Tuesday and then the final the week after LOL

OpenStudy (astrophysics):

That's brutallll

OpenStudy (anonymous):

Hopefully if I study hard enough for this test, I won't have to focus as much on it during finals week! X)

OpenStudy (astrophysics):

I hear ya, good luck! It's sort of insane how many methods we've learnt in such little time for DE's!

OpenStudy (anonymous):

Seriously! Haha. Best of luck to you too!

OpenStudy (irishboy123):

you drop Laplace Transforms pretty much as soon as you start questioning why it all works and i'm not sure that's a bad thing. cos they don't lead anywhere interesting.

OpenStudy (irishboy123):

fourier transforms are way more challenging, and way more esoteric too.

OpenStudy (astrophysics):

I hear that, use it in PDEs right

OpenStudy (anonymous):

lol, I don't think I'll have to learn Fourier for differential eq's.

OpenStudy (anonymous):

Well, ordinary

OpenStudy (astrophysics):

I might have to take PDE's :'(

OpenStudy (anonymous):

Ah =/ My major is Mechanical (probably changing to aerospace!) engineering. So I think this is my last pure-math class

OpenStudy (astrophysics):

But a pre-req for it here is real analysis...and I haven't done that

OpenStudy (astrophysics):

Ah, nice one